Eq. (3) : r=(mv)/(qB)

Aged Particles in Electromagnetic Fields Page 7 ditional Questions (can be answered outside of class) In the space below, give a derivation of Eq. (3). Hint: Apply Newtons second law to the centripetal direction: F)en cp
Answer #1

Mv^2 /r=Bqvsin theta

Because Magnetic force is source of centripital force while moving in circle angle between Vand B is 90 so

Mv^2/r=Bqv

Mv/r=Bq

So r =mv/Bq